OpenStudy (anonymous):

Find the sum of each geometric series 3-6+12... to 7 terms So I got S7=129 but have been told that is wrong. My work: S7= 3(1+2^7)/ 1+2 S7=129

OpenStudy (loser66):

I don't know why, I got that answer, too.

OpenStudy (anonymous):

Well, I guess they could be wrong too. But I tried doing it manually (the pattern being multiplied by -2) and I got 192.

OpenStudy (loser66):

nope, 192 is the 7th term

OpenStudy (loser66):

S7 is sum of 7 terms, not the 7th term

OpenStudy (loser66):

@jim_thompson5910 Please

OpenStudy (anonymous):

Oh, okay. So there is a difference? what would 7 terms be then?

OpenStudy (loser66):

you and me got S7 = 129 and you said that it's wrong, how dare I say something when I am wrong? wait for smarties.

OpenStudy (anonymous):

Hahaha, okay. Thanks for your time.

OpenStudy (loser66):

it's ok, friend, I want to study , too.

jimthompson5910 (jim_thompson5910):

first term: a = 3 common ratio: r = -2 Sequence is: 3, -6, 12, -24, 48, -96, 192, so the 7th term is 192

jimthompson5910 (jim_thompson5910):

and using the formula Sn = a*(1-r^n)/(1-r) gives you Sn = a*(1-r^n)/(1-r) S7 = 3*(1-(-2)^7)/(1-(-2)) S7 = 3*(1-(-128))/(1-(-2)) S7 = 3*(1+128)/(1+2) S7 = 3*(129)/(3) S7 = 129 So that confirms that the sum of the first 7 terms is 129

OpenStudy (anonymous):

Huh, okay. So my teacher was wrong then. Thank you!

jimthompson5910 (jim_thompson5910):

yeah it looks it, unless there's more to the story

OpenStudy (loser66):

glad to see that we are not wrong hehehe. thank you very much @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw